OpenStudy (anonymous):

If the simultaneous equations \[y= x^{2}-k\] y=x have only one solution, find k. A. -1 B. \[-\frac{ 1 }{ 4 }\] C. -4 D. \[\frac{ 1 }{ 4 }\] E. 1

hartnn (hartnn):

just put y=x in y=x^2-k what u get ?

OpenStudy (anonymous):

x=x^(2)-k k=x(x-1) what's the answer?

hartnn (hartnn):

the equation is x^2 -x-k=0 now u know how to find k if there is only one solution of quadratic equation ?

OpenStudy (anonymous):

just sub the answer into the equation? -.-

OpenStudy (anonymous):

is it B?

hartnn (hartnn):

yes, thats a way , but easier would be to equate discrimant to 0 know what discriminant is ?

OpenStudy (anonymous):

b^(2)-4ac (-1)^(2)-4(1)(-k) ?

hartnn (hartnn):

yes, equate that to 0

hartnn (hartnn):

1+4k=0 k= ?

OpenStudy (anonymous):

-1/4 thank you and I have many questions, would you kindly help me ?

hartnn (hartnn):

i will surely try...i am helping 2-3 ppl at a time...

OpenStudy (anonymous):

oh...sorry...I think I can ask another ppl... you are too busy :P

hartnn (hartnn):

open it as a new question...if not me, some1 will surely help you :)